OpenStudy (anonymous):

The circumference of a planet is about 26,000 miles. Find its diameter to the nearest mile.

OpenStudy (anonymous):

Circumference = (pi)(diameter). So substitute and solve.

OpenStudy (anonymous):

i keep working it out to be 81640? what am i doing wrong?

OpenStudy (anonymous):

what did you do to get 81640?

OpenStudy (anonymous):

took half of the circumference multiplyed by 2 then by pi

OpenStudy (anonymous):

how do i get the diameter? first?

OpenStudy (shamil98):

26000 = 3.14(d) so solve for d.

OpenStudy (anonymous):

That's where you went wrong. Circumference = (pi) x (diameter) 26,000 = (pi) x (diameter) To find diameter, divide both sides of ther equation by (pi), so you get diametr = 26,000/(pi) = 8,276 to the rearest mile.

OpenStudy (anonymous):

nearest*

OpenStudy (anonymous):

nice one thanks for clearing that up!